### v1

@@ 4 Experiments
A summary of our experimental results and comparison with previous methods is shown in Table 2. Below, we state the experimental setup and describe how different parts of the results are obtained and provide additional ablation study results.
@@ 4.1 Experimental Setup
We evaluate our methods on miniF2F (Zheng et al., 2021), a benchmark for formal Olympiad-level mathematics problems in four different formal languages. Specifically, we use the Isabelle/Isar part of the miniF2F-test dataset provided by Jiang et al. (2023), which provides an additional informal statement and informal draft for each problem. We use the PISA environment (Jiang et al., 2021) to interact with Isabelle 2022. Our <ATP>monospace-$<ATP>|$> method is the combination of 8 Isabelle proof methods (auto, simp, auto, blast, fastforce, eval, sos, arith, simp:field_simps, simp add:mod_simps) and Sledgehammer. Following Jiang et al. (2023), we set the timeout for any proof step and Sledgehammer as 10s and 120s, respectively. During verification, we run 12 PISA instances in parallel on one Intel(R) Xeon(R) Gold 6326 CPU @ 2.90GHz [^5] To reproduce our results, using the same CPU is a necessary requirement. See Appendix D for more discussion on the reproducibility issues..
@@
For few-shot prompting (and similarly for zero-shot prompting), we prepare few-shot examples that are the same with those of Jiang et al. (2023) except that the proof sketches are modified to new full proofs written by us. For $n$-shots prompting, we randomly sample $n$ examples from the candidates. By default, we use 1-shot prompting. See Appendix E for our specific prompt text, the reason of choosing 1-shot, and more details on the prompt construction process.
For multiple attempts of proof, we record the checked semi-proofs and skip them when obtaining them in the subsequent attempts. This will in fact harm the final performance since some tactics and Sledgehammer do not always give the same result under the same condition. Nevertheless, this practice can help us significantly reduce the elapsed time of the experiments.
@@ 4.2 Results of ProofAug
We do ablation study to validate the effectiveness of ProofAug (without the ERP module). We first get the DSP baseline results in our setting. Compared to the performance of the original DSP implementation on miniF2F-test, our DSP baseline is 9.9% higher (49.2% v.s 39.3% in 100 attempts) due to differences on proving environments [^6] See Appendix D for a discussion on the differences of theorem proving environment between ours and that of Jiang et al. (2023)., choice of prompting method and model difference. Then we compare ProofAug results with our DSP baselines. Under sample budgets of 1, 10 and 100, ProofAug achieves performance of 36.5%, 44.7% and 52.5%, significantly outperforming the baseline by 7.8%, 4.1% and 3.3%, respectively.
Note that ProofAug can also be applied to the proof sketches generated by DSP. Figure 3 shows how the number of proved theorems changes as the number of attempts increase w/ or w/o ProofAug for three prompting methods: DSP prompting, full proof prompting, and the more token-usage friendly zero-shot prompting. By w/o ProofAug, we mean the ‘Prove’ step in DSP, i.e., <ATP>monospace-$<ATP>|$> is tried at the sorrys in the proof sketch. It can be seen that while the performance of our few-shot prompting and zero-shot prompting w/o ProofAug is much lower than that of DSP prompting, ProofAug helps them achieve a comeback. This justifies our choice of generating full proofs instead of proof sketches.
@@ 4.3 Results of the ERP module
We try 500 attempts of proof for each problem, using zero-shot prompting for ProofAug w/ and w/o ERP module. When then ERP module is turned on, we record the cumulative number of queries and stop verifying when it reaches 500 for a fair comparison.
The result shows that ERP module can bring an extra 1.6% performance gain. If we do not restrict the number of queries, we can achieve a pass rate of 57.4%(+2.9% compared to the ProofAug without ERP).
@@ 4.4 Cumulative Results of Mixed Strategy Experiments
During the completion of this work, we have tried various setups to search for the best recipe. While some setups fail to achieve the best performance on the benchmark compared to results we present above, diverse ways of generating proofs help prove more theorems. The setups include: 1) Three versions of demonstration examples. for few-shot prompting generation and two versions of zero-shot prompt are used in our experiments. 2) Whether to include Informal drafts. The informal drafts could mislead LLMs. Although ProofAug can sometimes be a rescue, it is not always the case. Besides, we find some incorrect theorem statements in the miniF2F-test dataset during our experiments, so we build a curated dataset and part of the experiments are done on the curated version.
In total, we try each problem of the original miniF2F-test under a sample budget of around 1400 in total, and 2100 for the two versions in total. See Table 3 in Appendix F for a breakdown of these two numbers. On the original miniF2F-test, we solve 61.9% of the problems, which is 5.8% higher than the previous Isabelle SOTA. On the curated version, the solving rate comes to 66.0%, setting a new state of the art across all proof systems.
@@ 5 Discussion and Future Directions
Although this work demonstrates that our ProofAug method significantly improves the performance of the neural theorem prover on the miniF2F benchmark, there remains a notable gap in tackling IMO-level problems or research-level challenges within formal systems. How to incorporate ProofAug into the expert iteration methods (Xin et al., 2024a) or online learning methods (Lample et al., 2022) to further improve the performance is one direction of the future work.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Other Related Work
In the main body, Table 1 and Table 2 only include representative neural theorem proving methods that are most related to this work: they all target on solving Olympic-level mathematical problems with generative language models. Below we introduce other related work.
@@
Hierarchical Theorem Proving. We note that the idea of hierarchically proving theorems in declarative ITP systems has been explored in several previous works. Besides the proof-step based recursive proving method POETRY (Wang et al., 2024) discussed in Section 3.3, LEGO-Prover (Wang et al., 2023a) uses a skill library to record the previously proved lemmas and retrieves relative skills when instructing GPT series models to synthesize its proof based on a decomposed informal draft. It is a hierarchical proving method in the sense that the lemmas used for proving current theorem can be built on more low-level lemmas. From another point of view, LEGO-Prover uses all previous queries to solve the current theorem hierarchically. In contrast, Dong et al. (2024) does not record a skill library, but finishes the hierarchical proving from scratch. The key difference between these hierarchical proving methods and our ProofAug (without ERP module) is that we work in a bottom-up fashion to best exploit the power of traditional automation tools, contrary to their top-down frameworks that all require extra queries to the model when stepping into the next level proof. We argue that bottom-up fashion aligns more with the pre-trained corpus, thus making best use of the theorem proving knowledge possessed by the model. In contrast, top-down methods require in-context learning or fine-tuning to control the behavior of LMs, which could hurt the intrinsic ability of the model and bring extra complexity to the problem. This is likely the underlying reason why ProD-RL (Dong et al., 2024) fails to make significant improvements over the simple SFT method on miniF2F-test. Whereas our ProofAug serves as a play-and-plug module and can fall back to naive single-pass generation, being superior in efficiency, robustness and flexibility. Moreover, by adding the ERP module, we integrate the top-down and bottom-up fashions in a sample-efficient way.
@@
Declarative versus Procedural. The discussion of proof styles dates back to Harrison (1996). Wiedijk (2006) compares the proof styles of a wide range of different systems. Wiedijk (2012) discusses how to convert from procedural system to a declarative one. Avigad et al. (2021) introduces the two styles of proof in Lean and provides concrete examples.
@@
Other whole-proof generation works. Xin et al. (2024a); Wu et al. (2022) verify the effectiveness of expert iteration on neural theorem proving using the naive whole-proof generation method. They build the first-round theorem statements through autoformalization of the informal problems. First et al. (2023) employs Isabelle error message to enhance the direct full proof synthesis ability of LLM.
@@ Appendix B Examples of the DSP Issues
Figure 4 and Figure 5 show examples of the ‘Hard Conjecture’ issue and the ‘Complicated Draft’ issuse described in Section 1.
@@ Appendix C Discussion of Implementing ProofAug for Other Proof Assistants
Lean. Although Lean proofs can be completely tactic-based, block-structures starting with by are very common as well, especially for the proofs of Olympics-level math problems (refer to Xin et al. (2024b) for the example proofs they find). As a result, Lean proofs are also amendable for proof structure analysis with some modifications. As to the substitution of the <ATP>monospace-$<ATP>|$> method, firstly, there are built-in tactics such as simp, ring, nlinarith similar to the heuristics proof methods in Isabelle that have a predefined set of rules (facts) to use. Besides, there have been projects that can provide Lean tactics that discharge goals into SMT solvers already777lean-smt. It is hopeful that a Lean version of ProofAug can be constructed with the help of such tools by extra engineering efforts.
@@ Appendix D Theorem Proving Environment Setup Details and Discussion on the Reproducibility
Nevertheless, even using the same environment, it is still hard to produce exactly the same results due to the inherent randomness of these automation methods. The good news, however, is that for theorem proving, we are confident that we have made some progress when new theorems get proved, no matter whether this is completely reproducible or not. So we focus on making sure the obtained proofs are correct by verifying them again in Isabelle. We provide the found proofs in attachment to our code. Note that although there are cases where the proofs might fail due to timeout of the reconstructed smt or metis proofs, it does not mean that the theorems are not proved. It is just the limitation of Isabelle’s built-in tactics.
@@ Appendix G Additional Experimental Results
Appendix H An Example of How ProofAug Induce a Proof
Figure 9 shows an example of ProofAug inducing a proof from a failed proof proposal generated by the model. It can be seen that the initial proof fails inside the proofqed block of proving $a(2-a)\leq 1$ due to the wrong intermediate claim $2*a-a^{2}=(a-1)^{2}+1-a^{2}$. ProofAug finds that $a(2-a)\leq 1$ can be directly proved with the proof method $sos|$ given the previous proved fact $(a-1)^{2}\geq 0$.

### v2

@@ 4 Experiments
A summary of our experimental results and comparison with previous methods is shown in Table 2. Below, we state the experimental setup and describe how different parts of the results are obtained and provide additional ablation study results. For the results in Lean implementation, please refer to Appendix C.
4.1 Benchmark
We evaluate our methods on the Isabelle part of miniF2F (Zheng et al., 2021), a benchmark for formal Olympiad-level mathematics problems across four different formal languages. Unless otherwise specified, we use ‘miniF2F-test’ or ‘the original version of miniF2F-test’ or ‘the DSP-version miniF2F-test’ to refer to the dataset included in the code repository [^8] https://github.com/albertqjiang/draft_sketch_prove/blob/main/aligned_problems/new_complete_mini.jsonl of Jiang et al. (2023), which provides an additional informal statement and informal draft for each problem compared to the OpenAI miniF2F  [^9] https://github.com/openai/miniF2F/tree/main/isabelle. While both versions contain many mistakes and various extents of inaccuracies in the formal statements, previous work (Wang et al., 2023a; Zheng et al., 2024; Wang et al., 2024; Zhao et al., 2024) in general chooses to use this DSP-version for a fair comparison between different methods. Thus, in order to show the superiority of our ProofAug, we first follow previous work to do experiments on the DSP-version miniF2F-test.
However, as the performance of neural theorem provers has significantly improved recently, the ratio of unprovable theorems resulting from the mistakes to the total number of statements that provers fail to provide proofs becomes no longer ignorable, and is one of the possible reasons that Isabelle-based provers have fallen behind Lean-based ones on this cross-language benchmark. So, we build a curated dataset and part of our experiments are done on the curated version. [^10] You can find our curated version here: https://github.com/haoxiongliu/ProofAug/blob/main/isabelle_src/datasets/minif2f-test-curated.jsonl. See also Appendix H for the details of how we curate miniF2F.
@@ 4.2 Experimental Setup
The Isabelle version we use is Isabelle2022. We use the PISA environment (Jiang et al., 2021) to interact with Isabelle. Our <ATP>monospace-$<ATP>|$> method is the combination of 8 Isabelle proof methods (auto, simp, auto, blast, fastforce, eval, sos, arith, simp:field_simps, simp add:mod_simps) and Sledgehammer. Following Jiang et al. (2023), we set the timeout for any proof step and Sledgehammer as 10s and 120s, respectively. During verification, we run 12 PISA instances in parallel on one Intel(R) Xeon(R) Gold 6326 CPU @ 2.90GHz [^11] To reproduce our results, using the same CPU is a necessary requirement. See Appendix D for more discussion on the reproducibility issues..
@@
For few-shot prompting (and similarly for zero-shot prompting), we prepare few-shot examples that are the same as those of Jiang et al. (2023) except that the proof sketches are modified to new full proofs written by us. For $n$-shots prompting, we randomly sample $n$ examples from the candidates. By default, we use 1-shot prompting. See Appendix E for our specific prompt text, the reason for choosing 1-shot, and more details on the prompt construction process.
For multiple attempts of proof, we record the checked semi-proofs and skip them when obtaining them in the subsequent attempts. This, in fact, harms the final performance since some tactics and Sledgehammer do not always give the same result under the same conditions. Nevertheless, this practice can help us significantly reduce the elapsed time of the experiments.
@@ 4.3 Results of ProofAug
We do ablation studies to validate the effectiveness of ProofAug (without the ERP module). We first get the DSP baseline results in our setting. Compared to the performance of the original DSP implementation on miniF2F-test, our DSP baseline is 9.9% higher (49.2% v.s 39.3% in 100 attempts) due to differences in proving environments [^12] See Appendix D for a discussion on the differences of theorem proving environment between ours and that of Jiang et al. (2023)., choice of the prompting methods and model selection. Then we compare ProofAug results with our DSP baselines. Under sample budgets of 1, 10 and 100, ProofAug achieves performance of 36.5%, 44.7% and 52.5%, significantly outperforming the baseline by 7.8%, 4.1% and 3.3%, respectively.
Note that ProofAug can also be applied to the proof sketches generated by DSP. Figure 3 shows how the number of proved theorems changes as the number of attempts increases w/ or w/o ProofAug for three prompting methods: DSP prompting, full proof prompting, and the more token-usage friendly zero-shot prompting. By w/o ProofAug, we mean the ‘Prove’ step in DSP, i.e., <ATP>monospace-$<ATP>|$> is tried at the sorrys in the proof sketch. It can be seen that while the performance of our few-shot prompting and zero-shot prompting w/o ProofAug is much lower than that of DSP prompting, ProofAug helps them achieve a comeback. This justifies our choice of generating full proofs instead of proof sketches.
@@ 4.4 Results of the ERP module
We try 500 attempts of proof for each problem, using zero-shot prompting for ProofAug w/ and w/o ERP module. When the ERP module is turned on, we record the cumulative number of queries and stop verifying when it reaches 500 for a fair comparison.
The result shows that the ERP module can bring an extra 1.6% performance gain. If we do not restrict the number of queries, we can achieve a pass rate of 57.4%(+2.9% compared to the ProofAug without ERP).
@@ 4.5 Cumulative Results of Mixed Strategy Experiments
During the completion of this work, we have tried various setups to search for the best recipe. While some setups fail to achieve the best performance on the benchmark compared to results we present above, diverse ways of generating proofs help prove more theorems. The setups include: 1) Different versions of prompts. We have three versions of demonstration examples for few-shot prompt generation, and two versions of zero-shot prompt are used in our experiments. 2) Whether to include Informal drafts. The informal drafts could mislead LLMs. Although ProofAug can sometimes be a rescue, it is not always the case.
In total, we try each problem of the original miniF2F-test under a sample budget of around 1400 in total, and 2100 for the two versions in total. See Table 3 in Appendix F for a breakdown of these two numbers. On the original miniF2F-test, we solve 61.9% of the problems, which is 5.8% higher than the previous Isabelle SOTA. On the curated version, the solving rate comes to 66.0%.
@@ 5 Discussion and Future Directions
Although this work demonstrates that our ProofAug method significantly improves the performance of the neural theorem prover on the miniF2F benchmark, there remains a notable gap in tackling IMO-level problems or research-level challenges within formal systems. How to incorporate ProofAug into expert iteration methods (Xin et al., 2024a) or online learning methods (Lample et al., 2022) to further improve the performance is one direction of future work.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Other Related Work
In the main body, Table 1 and Table 2 only include representative neural theorem proving methods that are most related to this work: they all focus on solving Olympic-level mathematical problems with generative language models. Below we introduce other related work.
@@
Hierarchical Theorem Proving. We note that the idea of hierarchically proving theorems in declarative ITP systems has been explored in several previous works. Besides the proof-step based recursive proving method POETRY (Wang et al., 2024) discussed in Section 3.3, LEGO-Prover (Wang et al., 2023a) uses a skill library to record the previously proved lemmas and retrieves relative skills when instructing GPT series models to synthesize its proof based on a decomposed informal draft. It is a hierarchical proving method in the sense that the lemmas used for proving the current theorem can be built on more low-level lemmas. From another point of view, LEGO-Prover uses all previous queries to solve the current theorem hierarchically. In contrast, Dong et al. (2024) does not record a skill library, but finishes the hierarchical proving from scratch. The key difference between these hierarchical proving methods and our ProofAug (without ERP module) is that we work in a bottom-up fashion to best exploit the power of traditional automation tools, contrary to their top-down frameworks that all require extra queries to the model when stepping into the next level proof. We argue that bottom-up fashion aligns more with the pre-trained corpus, thus making best use of the theorem proving knowledge possessed by the model. In contrast, top-down methods require in-context learning or fine-tuning to control the behavior of LMs, which could hurt the intrinsic ability of the model and bring extra complexity to the problem. This is likely the underlying reason why ProD-RL (Dong et al., 2024) fails to make significant improvements over the simple SFT method on miniF2F-test. Whereas our ProofAug serves as a play-and-plug module and can fall back to naive single-pass generation, being superior in efficiency, robustness, and flexibility. Moreover, by adding the ERP module, we integrate the top-down and bottom-up fashions in a sample-efficient way.
While finalizing this paper, the authors became aware of a concurrent and independent work Lu et al. (2024), which proposes PALM, that shares similar high-level ideas with ProofAug. However, our work is distinct in several key aspects: 1) PALM works in Coq (Barras et al., 1999), a procedural proof language where declarative modes (such as the ‘Mathematical Proof Language’ introduced in the Coq reference manual) are less common. Since tactics in the procedural style proof languages are designed in the spirit that the user decides what tactic to use next by observing the resulting goal or proof state after using the last tactic, it is often the case that new variables or facts are introduced by one tactic and the subsequent tactics need to specify the names of these facts. Thus, performing whole-proof generation as PALM does for Coq is generally harder than in the declarative style proofs and could be detrimental to the performance. Besides, in Coq proofs, there are few explicit intermediate conjectures or block structures as in those in Isar or Lean 4. Thus, for each tactic, the backtracking procedure of PALM simply goes back to the proof state before the unrepairable error and continues backtracking, except for the bullet point structure that is equivalent to the ‘case’ structure in Isabelle/Isar (rather than the more common ‘have’ method). Consequently, the proofs found by PALM mostly show scarce high-level structures as we can observe from the results provided in its code repository  [^13] https://github.com/lachinygair/PALM/tree/main/evaluation/proof. In contrast, our ProofAug departs from the DSP framework and takes an ‘augmentation’ perspective to solve the issues of DSP. We first determine all compatible semi-proofs, then sequentially prune the semi-proofs that become impossible to be extended into a full proof after each time we apply <ATP>monospace-$<ATP|>$ to an intermediate conjecture. This procedure makes ProofAug fit with the ‘sketch’ procedure of DSP that translates the informal proof into intermediate conjectures, and each semi-proof is expected to align with the corresponding part of the informal proof. Note that such correspondence is hard to achieve for procedural proofs and obligating the LLMs to use the uncommon declarative mode or ‘assert’ keywords in Coq could be detrimental to the performance. Although in Appendix C we also discuss how to transform a procedural style proof language into a declarative one in the spirit described in Wiedijk (2012) and then train models and apply ProofAug accordingly, this procedure needs to fine-tune a model on enough data, which is in general expensive. In total, ProofAug is the choice for theorem proving models such as the DeepSeek-Prover series and the Kimina-Prover (Wang et al., 2025) series that achieve SOTA performance on the cross-language benchmark miniF2F, and provide detailed solutions and code. 2) Our work further explores the Efficient Recursive Proving module and shows it can boost the performance with a relatively small number of queries. In fact, the counterpart of the ERP module for procedural-style proofs should be some kind of truncate-and-resume tree-search algorithm (that only adds new nodes based on the initial proof proposal) proposed in Xin et al. (2024b). However, since the proof style adopted by Deepseek-Prover-V1.5 series models is rather declarative, RMaxTS becomes inefficient for them due to the potential reasons we analyzed in Section 3.3. 3) We make extra contributions such as exploring extended mixture of strategies to enhance the performance, curating the Isabelle version of miniF2F-test and finally achieving a cumulative pass rate of 66.0% on this benchmark, while the performance of Coq-based methods remains unclear on the problems of miniF2F since the Coq version is not included by now.
@@
Declarative versus Procedural. The discussion of proof styles dates back to Harrison (1996). Wiedijk (2006) compares the proof styles of a wide range of different systems. Wiedijk (2012) discusses how to convert from a procedural system to a declarative one. Avigad et al. (2021) introduces the two styles of proof in Lean and provides concrete examples.
@@
Other whole-proof generation works. Xin et al. (2024a); Wu et al. (2022) verify the effectiveness of expert iteration on neural theorem proving using the naive whole-proof generation method. They build the first-round theorem statements through autoformalization of the informal problems. First et al. (2023) employs Isabelle error messages to enhance the direct full proof synthesis ability of LLM.
@@ Appendix B Examples of the DSP Issues
Figure 4 and Figure 5 show examples of the ‘Hard Conjecture’ issue and the ‘Complicated Draft’ issue described in Section 1.
@@ Appendix C Discussion of Implementing ProofAug for Other Proof Assistants
Lean. Although Lean proofs can be completely tactic-based, block-structures starting with by are very common as well, especially for the proofs of Olympics-level math problems (refer to Xin et al. (2024b); Wang et al. (2025) for the example proofs they find). As a result, Lean proofs are also amendable for proof structure analysis, so we also build a simple implementation of ProofAug in Lean. As to the substitution of the <ATP>monospace-$<ATP>|$> method, we use the combination of Aesop(Limperg & From, 2023), omega and our hand-made heuristic try norm_num [*]; try field_simp [*] at *; try ring_nf at *; try nlinarith. For the proof structure analysis process, given a line of code, we manually infer whether it is intended to be part of a statement or tactic from its indent and context. Our experiment result shows that ProofAug improves the performance of Kimina-Prover-Preview-Distill-1.5B from 44.3% to 50.4% on miniF2F. The improvement can be further enhanced if we can obtain a more proper control of the parsing and structure analysis process and get better substitutions of the <ATP>monospace-$<ATP>|$> method. Readers may refer to our code for more implementation details.
@@ Appendix D Theorem Proving Environment Setup Details and Discussion on the Reproducibility
Nevertheless, even using the same environment, it is still hard to produce exactly the same results due to the inherent randomness of these automation methods. The good news, however, is that for theorem proving, we are confident that we have made some progress when new theorems get proved, no matter whether this is completely reproducible or not. So we focus on making sure the obtained proofs are correct by verifying them again in Isabelle. We provide the found proofs in attachment to our code. Note that although there are cases where the reconstructed proof fails due to the timeout of Isabelle built-in methods such as smt and metis, it does not mean that the theorems are not proved. It is only the limitation of these tactics.
@@ Appendix G Additional Experimental Results
Appendix H Curation of the Isabelle part of miniF2F
Our curation of DSP-version miniF2F test split mainly includes two types: 1) Typo fix: Statements with typos in variable names or numbers are mostly completely wrong and unprovable. 2) Modifying the type of the variables from nat to int for theorems that could potentially involve the minus operation in the proofs. This is important because, for example, the operation (2::nat) - (3::nat) is valid and expected to be (0::nat) in Isabelle/HOL. Although sometimes the original statement can still be proved in the sense that the resulting 0 also satisfies the constraints in the statement, it does not express the intent of the original problem. Figure 9 shows such an example.
After finishing the curation process, the authors get to learn about that there is a FAIR version repository of miniF2F [^14] https://github.com/facebookresearch/miniF2F that is actively in maintenance and most of the typos and inappropriate uses of minus of nat type numbers we find in the DSP-version have been fixed in the latest commits. Nevertheless, several typos and some inappropriate use of nat type remain. Thus, we still make a pull request [^15] https://github.com/facebookresearch/miniF2F/pull/18 to the FAIR miniF2F repository to update part of our curation to the upstream.
Differences:
theorem imo_2001_p6:
- fixes a b c d ::nat
+ fixes a b c d ::int
assumes "0 < a \<and> 0 < b \<and> 0 < c \<and> 0 < d"
and "d < c"
and "c < b"
and "b < a"
- and "a * c + b * d = (b + d + a - c) * (b + d - a + c)"
+ and "a * c + b * d = (a + b - c + d) * (- a + b + c + d)"
shows "\<not> prime (a * b + c * d)"
Figure 9: How we fix the imo_2001_p6 in DSP-version miniF2F-test. The "a * c + b * d = (b + d + a - c) * (b + d - a + c)" constraint does not express what it intends in the original statement when $a>b+d$.
Appendix I An Example of How ProofAug Induce a Proof
Figure 10 shows an example of ProofAug inducing a proof from a failed proof proposal generated by the model.